Lake Charles Seafood makes 450 wooden packing boxes for f fresh seafood per day, working in two 10-hour shifts. Due to increased demand, plant managers have decided to operate three 8-hour shifts per day instead. The plant is now able to produce 650 boxes per day a) Before the chan After the change, the new productivity level per day 27.08 boxes/hour (round your response to two decimal places) b) Based on the changes made the percent increase in productivity 20 3 % enter your response as a percentage rounded to two decimal places c) If production is increased to 800 boxes per day (with the three 8-hour shifts), the new productivity- 33 33 boxes/hour (round your response to two decimal places) nge in work rules, the companys productivity per day 2250 boxeshour (round your response to two decimal places)
Answer #1

Please provide rating..

Answer a)
Productivity = Unit produced / Hr worked
Before After
i Unit produced 450 650
ii Hr worked =10*2 20 =8*3 24
iii=i/ii Productivity       22.50       27.08
Answer b)
% increase in productivity = (27.08-22.5)/22.5 20.36%
Answer c)
i Unit produced = 800
ii Hr wroked= 24
iii=i/ii New productivity =       33.33